Came this morning to find that all the BHD optics watchdogs tripped. Watchdogs were restored but the all ADCs were stuck.
We realized that the C1SU2 model crashed.
We run /scripts/cds/restartAllModels.sh to reboot all machines. All the machines rebooted and burt restored to yesterday around 5 PM.
The optics seem to have returned to good alignment and are damped.